# OpenFGL: A Comprehensive Benchmarks for Federated Graph Learning

### v1

@@ A.4. Robustness Simulation Description
Given the practical applications driving FGL studies, the pivotal goal of various FGL approaches should be their effective deployment in real-world industrial scenarios. Consequently, conducting a thorough evaluation of the robustness of existing methods becomes essential. In our proposal, besides exploring the generalization of current methods across various federated data simulation settings as discussed in Sec. 3.1, we draw insights from common business challenges encountered in industrial scenarios. Specifically, we additionally integrate the following experimental setups to provide a comprehensive evaluation for industrial research projects from a robustness perspective.
@@ A.7. Hyperparameter Settings
Regarding graph-specific data simulation strategies in Table 5, to enhance readability and avoid complex figures or tables, we default to using 10-client label Dirichlet (i.e., Label Distribution Skew and $\alpha=1$) and Metis partitioning (i.e., Metis-based Community Split) separately for the Graph-FL and Subgraph-FL scenarios. The former is inspired by data Non-iid simulation in CV (Li et al., 2020; Karimireddy et al., 2020; Li et al., 2021), while the latter is inspired by prevalent data simulation strategies in current FGL studies (Baek et al., 2023; Li et al., 2023b, 2024b). Experimental evaluations of the generalization of existing methods across different data simulation scenarios can be found in Sec. 4.2. Furthermore, in the selection of local backbone models for graph learning, we choose prevalent GIN and GCN models, applied to Graph-FL and Subgraph-FL, respectively. Notably, we experiment with multiple datasets and baselines in separate modules and use graph/node classification to report experimental results to further avoid complex charts, making the results more reader-friendly.
@@ A.8.3. Technique Details of FedDEP
In addition, we have implemented the FedDEP (Zhang et al., 2024c) algorithm within the OpenFGL framework. FedDEP achieves noise-free edge-level DP by employing random sampling, ensuring robust privacy preservation without compromising data integrity. To tackle the issue of cross-subgraph missing neighbors, FedDEP incorporates an advanced deep neighbor generation module known as DGen, which enhances the model’s capability to generate and integrate missing neighbors effectively. To further optimize computational efficiency within each client, local GNN embeddings are clustered to create sets of missing neighbor prototypes. Moreover, to reduce inter-client communication overhead, FedDEP introduces a pseudo-federated learning approach, where these prototype embeddings are shared across the system before the training of DGen, thereby streamlining the collaborative learning process while maintaining model accuracy and privacy.
@@ A.8.4. Future Research Directions
This section introduces several promising research directions in DP-based FGL. Currently, there are few privacy-preserving algorithms specifically tailored for FGL. Given the distinctive properties of graph data, traditional privacy mechanisms commonly employed in FL fail to seamlessly extend to this domain. Therefore, it is crucial to develop new mechanisms for FGL that should consider the unique characteristics of GNNs to achieve an optimal trade-off between privacy protection and model utility. Furthermore, the convergence analysis of differentially private FGL algorithms is still an open research problem. A tight convergence upper bound not only theoretically assures rapid convergence, but also facilitates an empirical examination of how various hyperparameters influence convergence rates. Such insights are pivotal for fine-tuning parameters or the development of innovative optimization strategies (Li and Lyu, 2024). Existing analyses, such as those presented in (Wei et al., 2020; Li et al., 2022b), do not consider GNN-specific processes like propagation and aggregation. Consequently, there is an urgent need for advanced convergence analysis approaches suitable for DP-FGL. Another important research direction is to evaluate the performance of differentially private FGL algorithms against malicious attacks, such as inference attacks (Nasr et al., 2019; Luo et al., 2021) and reconstruction attacks (Zhu et al., 2019; Geiping et al., 2020). This research will contribute significantly to the robustness and reliability of DP-based FL.
@@ A.9. Federated Heterogeneous Graph Learning
Relation FedDA (Gu et al., 2023) Simple-HGN (Lv et al., 2021) DBLP Dominant
@@ A.9.1. Relation Type Sharing
Compared to FGL algorithms, FedDA (Gu et al., 2023) analyzes the unique characteristics of FL on HGs: that is, only a small amount of client parameters need to be uploaded in each communication round to achieve rapid convergence. Based on these findings, FedDA proposes a dynamic activation strategy, which achieves efficient training by dynamically selecting a subset of clients for each round of aggregation. Furthermore, considering the complex relations in HGs, FedDA introduces the D-HGN model, which decouples the parameters of relation types to allow partial updates of relation parameters on the server side instead of the entire model parameters.
@@ A.9.3. Basic FHGL on OpenFGL
Although in FHGL, different HGNNs (Schlichtkrull et al., 2018; Wang et al., 2019; Hu et al., 2020a) and FL methods (McMahan et al., 2017; Gao et al., 2022; Li et al., 2021; Karimireddy et al., 2020) can be combined to achieve the basic framework, it lacks distributed characteristics. For example, in centralized learning, the attention mechanism is considered an effective method of identifying relations (Lv et al., 2021). However, in Table 13, HAN and HGT based on different attention mechanisms perform worse in the FHGL scenario compared to the RGCN. We speculate that this is because the attention mechanism requires additional parameters, making it more prone to over-fitting and stronger local biases. Additionally, the information loss caused by meta-paths is further amplified in FGL, resulting in HAN achieving the worst performance. It is worth mentioning that in most cases, the performance of the same HGNN model shows almost no significant differences across different FL methods. Therefore, for distributed HGs, designing more reasonable HGNNs seems to be a more effective approach. OpenFGL has user-friendly extensibility, allowing users to quickly experiment with their own HGNN and FHGL strategies. Given the practical value of FL and the more realistic modeling scenarios of HG, OpenFGL will inspire more users to conduct research on FHGL.

### v2

@@ A.4. Robustness Simulation Description
Given the practical applications driving FGL studies, the pivotal goal of various FGL approaches should be their effective deployment in real-world industrial scenarios. Consequently, conducting a thorough evaluation of the robustness of existing methods becomes essential. In our proposal, besides exploring the generalization of current methods across various federated data simulation settings as discussed in Sec. LABEL:sec:benchmark_scenario, we draw insights from common business challenges encountered in industrial scenarios. Specifically, we additionally integrate the following experimental setups to provide a comprehensive evaluation for industrial research projects from a robustness perspective.
@@ A.7. Hyperparameter Settings
Regarding graph-specific data simulation strategies in Table 5, to enhance readability and avoid complex figures or tables, we default to using 10-client label Dirichlet (i.e., Label Distribution Skew and $\alpha=1$) and Metis partitioning (i.e., Metis-based Community Split) separately for the Graph-FL and Subgraph-FL scenarios. The former is inspired by data Non-iid simulation in CV (Li et al., 2020b; Karimireddy et al., 2020; Li et al., 2021a), while the latter is inspired by prevalent data simulation strategies in current FGL studies (Baek et al., 2023; Li et al., 2023b, 2024e). Experimental evaluations of the generalization of existing methods across different data simulation scenarios can be found in Sec. 4.2. Furthermore, in the selection of local backbone models for graph learning, we choose prevalent GIN and GCN models, applied to Graph-FL and Subgraph-FL, respectively. Notably, we experiment with multiple datasets and baselines in separate modules and use graph/node classification to report experimental results to further avoid complex charts, making the results more reader-friendly.
@@ A.8.3. Technique Details of FedDEP
In addition, we have implemented the FedDEP (Zhang et al., 2024d) algorithm within the OpenFGL framework. FedDEP achieves noise-free edge-level DP by employing random sampling, ensuring robust privacy preservation without compromising data integrity. To tackle the issue of cross-subgraph missing neighbors, FedDEP incorporates an advanced deep neighbor generation module known as DGen, which enhances the model’s capability to generate and integrate missing neighbors effectively. To further optimize computational efficiency within each client, local GNN embeddings are clustered to create sets of missing neighbor prototypes. Moreover, to reduce inter-client communication overhead, FedDEP introduces a pseudo-federated learning approach, where these prototype embeddings are shared across the system before the training of DGen, thereby streamlining the collaborative learning process while maintaining model accuracy and privacy.
@@ A.8.4. Future Research Directions
This section introduces several promising research directions in DP-based FGL. Currently, there are few privacy-preserving algorithms specifically tailored for FGL. Given the distinctive properties of graph data, traditional privacy mechanisms commonly employed in FL fail to seamlessly extend to this domain. Therefore, it is crucial to develop new mechanisms for FGL that should consider the unique characteristics of GNNs to achieve an optimal trade-off between privacy protection and model utility. Furthermore, the convergence analysis of differentially private FGL algorithms is still an open research problem. A tight convergence upper bound not only theoretically assures rapid convergence, but also facilitates an empirical examination of how various hyperparameters influence convergence rates. Such insights are pivotal for fine-tuning parameters or the development of innovative optimization strategies (Li and Lyu, 2024). Existing analyses, such as those presented in (Wei et al., 2020; Li et al., 2022c), do not consider GNN-specific processes like propagation and aggregation. Consequently, there is an urgent need for advanced convergence analysis approaches suitable for DP-FGL. Another important research direction is to evaluate the performance of differentially private FGL algorithms against malicious attacks, such as inference attacks (Nasr et al., 2019; Luo et al., 2021) and reconstruction attacks (Zhu et al., 2019; Geiping et al., 2020). This research will contribute significantly to the robustness and reliability of DP-based FL.
@@ A.9. Federated Heterogeneous Graph Learning
Relation FedDA (Gu et al., 2023b) Simple-HGN (Lv et al., 2021) DBLP Dominant
@@ A.9.1. Relation Type Sharing
Compared to FGL algorithms, FedDA (Gu et al., 2023b) analyzes the unique characteristics of FL on HGs: that is, only a small amount of client parameters need to be uploaded in each communication round to achieve rapid convergence. Based on these findings, FedDA proposes a dynamic activation strategy, which achieves efficient training by dynamically selecting a subset of clients for each round of aggregation. Furthermore, considering the complex relations in HGs, FedDA introduces the D-HGN model, which decouples the parameters of relation types to allow partial updates of relation parameters on the server side instead of the entire model parameters.
@@ A.9.3. Basic FHGL on OpenFGL
Although in FHGL, different HGNNs (Schlichtkrull et al., 2018; Wang et al., 2019; Hu et al., 2020a) and FL methods (McMahan et al., 2017; Gao et al., 2022; Li et al., 2021a; Karimireddy et al., 2020) can be combined to achieve the basic framework, it lacks distributed characteristics. For example, in centralized learning, the attention mechanism is considered an effective method of identifying relations (Lv et al., 2021). However, in Table 16, HAN and HGT based on different attention mechanisms perform worse in the FHGL scenario compared to the RGCN. We speculate that this is because the attention mechanism requires additional parameters, making it more prone to over-fitting and stronger local biases. Additionally, the information loss caused by meta-paths is further amplified in FGL, resulting in HAN achieving the worst performance. It is worth mentioning that in most cases, the performance of the same HGNN model shows almost no significant differences across different FL methods. Therefore, for distributed HGs, designing more reasonable HGNNs seems to be a more effective approach. OpenFGL has user-friendly extensibility, allowing users to quickly experiment with their own HGNN and FHGL strategies. Given the practical value of FL and the more realistic modeling scenarios of HG, OpenFGL will inspire more users to conduct research on FHGL.